Accelerator Update for March 18, 2009 - March 20, 2009 |
Wednesday March 18 The day shift began with Operations monitoring the Tevatron (TeV) store 6903 (72.02E30) and the Recycler stash (236.1E10), with the Antiproton Source stacking (31.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 9:11 AM, a Linac quadrupole power supply (QPS113) tripped its breaker. It reset. At 9:31 AM, a Linac Klystron RF station (KRF6) tripped off on a LCW flow fault indication. It reset. TeV experts began and end-of-store study at 9:56 AM. At 11:39 AM, Operations reported a flurry of MI beam aborts at injection. MI and Booster experts investigated and made adjustments. TeV experts completed their study an Operations terminated store 6903 at 11:42 AM. The Neutron Therapy Facility (NTF) reported at 12:03 PM that they had treated their last patient for the day. Operations prepared D0 and CDF for access at 12:10 PM. Operations prepared the TeV sectors A-E for access at 12:27 PM. CDF came out of access at 1:39 PM. At 1:50 PM, BRF15 tripped off for the first of eight times in a little over an hour. It reset. Vacuum experts reported at 2:05 PM that they had completed their work on the TeV sector A3 roughing pump. Interlock experts reported at 2:49 PM that they were beginning their safety system tests on the A1 and P1 transfer lines. The tests were completed in about 30 minutes. At 3:29 PM, MIRF3 began tripping off. An RF expert began repairs and had it back online within 30 minutes. Operations established beam to the Meson MTest experiment at 4:01 PM. Operations reported at 4:05 PM that the Recycler had stashed antiprotons seven times during the day shift. At 4:26 PM, FESS personnel reported that they found a chilled water leak in Booster. The pictures below show the leak and their temporary repair.
Operations put the TeV through a dry squeeze at 4:46 PM. TeV experts began a study period at 5:21 PM. Operations disabled the MTest beam at 6:07 PM. Operations put the TeV through a wet squeeze at 6:21 PM. At 7:12 PM, Operations reported that the TeV sector D1 BPMs had caused an error in reading the sequencer. Operations put the TeV into shot setup at 7:15 PM. At 8:29 PM, Operations reported more problems with the TeV sector D1 BPM house. Operations established store 6906 at 8:36 PM with an initial luminosity of 314.49E30. At 10:07 PM, Operations reported that the TeV tunes were not reading back properly. An operator restarted the DCE04 machine. Operations reported at midnight that the Recycler had stashed three times during the evening shift. Thursday March 19 The midnight shift began with Operations monitoring store 6906 (177.88E30) and the stash (182.0E10), with Pbar stacking (14.9E10), and with NuMI and MiniBooNE taking beam. At 12:20 AM, LRF3 tripped off. It reset. At 12:39 AM, BRF15 tripped off. It reset At 12:52 AM, a Booster vacuum pump (IP61) tripped off. It reset. At 12:58 AM, BRF4 and 15 tripped off. They reset. At 1:01 AM, BRF4 tripped off for the first of two times in nine minutes and BRF15 tripped off for the first of seven times in five hours and 30 minutes. They reset. Operations called in a Booster expert at 2 AM to help with all the RF trips. Operations reported at 8 AM that the Recycler had stashed nine times during the midnight shift. Operations terminated store 6906 at 9:18 AM. Operations put the TeV into shot setup at 9:30 AM. At 9:57 AM, Operations reported regulation problems with a Linac quadrupole power supply (L:QTM3E). At 11:03 AM, a Booster power supply (B:MP03) tripped off on a “smoke detector” indication. Experts investigated and reported that its choke had burned up. They began repairs. Operations put Pbar into standby mode at 11:14 AM due to the MP03 repairs. Operations established store 6907 at 11:15 AM with an initial luminosity of 301.68E30. A Linac expert reported at 11:33 AM that he would be working on KRF5 and 6. A Linac expert reported at 11:37 AM that he would be working on QPS101, 102, and 117. Operations prepared Booster for access at 11:43 AM. FESS personnel reported at 1:14 PM that they were beginning work on the Pbar AP10 air compressor. Linac experts reported at 1:52 PM that they were running study beam pulses to the dump. At 3:19 PM, the NuMI beam permit tripped off when a power supply’s (E:V100) output drifted low. Everything reset. Booster came out of access at 3:29 PM. Experts had completed their work on MP03. Operations reported at 4:02 PM that the Recycler had stashed once during the day shift. Operations resumed sending beam to NuMI and Pbar at 4:09 PM. Operations established beam to MTest at 4:46 PM. A MI expert reported at 5:54 PM that she had checked coalescing. She also reported that the damper system was rebooted, but the reboot did not help the problem with bad beam positions. She will continue her investigations on Friday. A damper expert reported that he had replaced an amplifier, which should improve the proton quality. Operations disabled the MTest beam at 6:15 PM. At 6:46 PM, the P1 transfer line beam permit tripped off when power supply I:V714 drifted low. Everything reset. At 8:07 PM, TRF2 tripped off. This caused a small loss of luminosity. Operations reported at 8:23 PM that operators had completed a 750 KeV area tune. At 10:29 PM, LRF3 tripped off. It reset. Operations reported at midnight that the Recycler had stashed six times during the evening shift. Friday March 20 The midnight shift began with Operations monitoring store 6907 (65.42E30) and the stash (219.6E10), with Pbar stacking (18.3E10), and with NuMI and MiniBooNE taking beam. At 2:13 AM, CDF experimenters reported that they were having problems accessing their electronic logbooks. They were still taking data. Operations terminated store 6907 at 6:38 AM. Operators reported at 7:01 AM that TRF2 would not reset. They contacted an RF expert. Water experts reported at 7:19 AM that they would begin filling the TeV LCW system. NTF reported at 7:23 AM that they began treatment on their first of three patients for today. Operations reported at 8:22 AM that the Recycler had stashed nine times during the midnight shift. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/20/2009 email Fermilab |
| Security, Privacy, Legal | |